OpenStudy (anonymous):

Find the exact value of sin(arctan(2)). For full credit, explain your reasoning.

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle \(\theta\) with \(\tan(\theta)=2\)

OpenStudy (anonymous):

in other words \(\theta=\arctan(2)\) and you want \(\sin(\theta)\) all that is missing is the hypotenuse

OpenStudy (anonymous):

which you do in your head via pythatoras \[h=\sqrt{1^2+2^2}=\sqrt{5}\]

OpenStudy (anonymous):

now it should be easy to find \(\sin(\theta)=\sin(\arctan(2))\) it is opposite over hypotenuse

OpenStudy (anonymous):

How do you know the adjacent side is 1?

OpenStudy (anonymous):

i made it up \[\tan(\theta)=2\] so the ration of the "opposite" side to the "adjacent" sides is 2

OpenStudy (anonymous):

it is easiest to make the opposite side 2 and the adjacent side 1, but i could have made the opposite side 200 and the adjacent side 100 or opposite side 12 and the adjacent side 6

OpenStudy (anonymous):

all the ratios would work out the same, just easiest to make it 2 and 1